If A and B are events with P(A)=0.5, P(B)=0.3, P(A OR B)=0.67, find P(A AND B).
Answer #1

P(A) = 0.5

P(B) = 0.3

P(A or B) = 0.67

P(A or B) = P(A) + P(B) - P(A and B)

0.67 = 0.5 + 0.3 - P(A and B)

P(A and B) = 0.8 - 0.67

P(A and B) = 0.13